OpenStudy (anonymous):

what is the angle between -2.5+j4.33 and 4.33-j2.5??plz help..

OpenStudy (harsimran_hs4):

use cross product or dot product

OpenStudy (harsimran_hs4):

did you manage to work out something??

OpenStudy (anonymous):

(-2.5+j4.33)(4.33-j2.5) like this??

OpenStudy (harsimran_hs4):

take any dot product or cross ......do you know how to take these?

OpenStudy (anonymous):

can you write how to do it?

OpenStudy (harsimran_hs4):

ok 1. dot product is \[a . b = \left| a \right| \left| b \right| \cos \theta \] where theta is the angle b/w the two vectors |a| = magnitude of the vector a

OpenStudy (harsimran_hs4):

lets do it step by step 1. find the magnitude of both the vectors and tell me

OpenStudy (anonymous):

24.9989 for the both vectors..

OpenStudy (harsimran_hs4):

you also need to take squareroot of the expression for magnitude..isn`t it?

OpenStudy (anonymous):

* sq. root 24.9989

OpenStudy (harsimran_hs4):

cool now plug in the values in the formula i gave you above and tell me

OpenStudy (anonymous):

\[ \left|\sqrt{24.9989} \right| \times \left|\sqrt{24.9989} \right| \cos \theta\]

OpenStudy (anonymous):

like that??

OpenStudy (harsimran_hs4):

yes!! what about LHS

OpenStudy (anonymous):

150?? that correct?

OpenStudy (harsimran_hs4):

21.65 i suppose it`s just simple multiplication of i terms and j terms and then their addition

OpenStudy (anonymous):

how did you get 21.65?

OpenStudy (harsimran_hs4):

-2.5 * 4.33 + 4.33 * -2.5 -21.65

OpenStudy (anonymous):

150 is an angle i mean...sorry..

OpenStudy (harsimran_hs4):

i a hurry !! YES answer is correct Cool

OpenStudy (anonymous):

can you help me to express this sq. root of 15 + j 20??

OpenStudy (harsimran_hs4):

yes just give me a moment

OpenStudy (harsimran_hs4):

ok consider 5(3 + j 4) = 5( 4 -1 + i 2 *2) = 5(2^2 + i^2 + 2*2*i) = 5 * (2 + i)^2 so squareroot is \[\sqrt{5}(2 + i)\]

OpenStudy (harsimran_hs4):

did you get it or any problem?

OpenStudy (anonymous):

tnx..

OpenStudy (harsimran_hs4):

:)